OpenStudy (anonymous):

What is cos 2x when tan= -3/4 ?

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\] what you need to answer is "if \(\tan(x)=-\frac{3}{4}\) then what is \(\cos(x)\) ?"

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

yes indeed your only job now is to compute \[2\times(\frac{4}{5})^2-1\]

OpenStudy (anonymous):

33/25 is that the answer ?

OpenStudy (anonymous):

Is the answer 7/25?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to the second one, not the first one

OpenStudy (anonymous):

One more question what is the sin 2x when tan = -3/4 2/1 x -3/5 x 4/5 = -24/25 I got that is it right?